using the ratios 24/32 = 6/(x -1) or 24(x -1) =32 x 6 solve for x
Can you please explain so I can do the other questions thank you.

Answer by mananth(16946)

24(x -1) =32 *6
24x-24 =192
+24
24x=192+24
24x=216
/24
x=9